OpenStudy (anonymous):

When the smaller of two consecutive integers is added to four times the larger, the result is 49. Find the integers.

OpenStudy (anonymous):

i guess 10 and 11 lets check \[10+4\times 11]10+44=54\] hope i guessed too big try 8 and 9

OpenStudy (anonymous):

How did you get 8 and 9

OpenStudy (anonymous):

i didn't i just guessed still didn't check it is \[5+4\times 9=49\]?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok that was a typo i meant is \[8+4\times 9=49\]

OpenStudy (anonymous):

the answer is no it is not, but \[9+4\times 10=49\] so the numbers are 9 and 10

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

Thanks for your help bro

OpenStudy (anonymous):

yw your math teacher wants you to write \[x+4(x+1)=49\] and solve that you will get \(x=9\)

OpenStudy (anonymous):

Yeah I figure it out when I looked at an example